OpenStudy (anonymous):

compute the fifth degree Taylor polynomial of the function f(x)=cos(0.3x) at a=0 and Find the largest integer k such that for all x for which |x|<1 , the Taylor polynomial T 5 (x) approximates f(x) with error less than 1/10^k .

OpenStudy (anonymous):

i already found the fifth degree taylor poly. which is 1-.3^2/2(x^2)+(.3^4)/24*(x^4) im just not sure how to calculate the second part

OpenStudy (anonymous):

see this

OpenStudy (anonymous):

thanks but im not sure what im suppose to be seeing in the graph? sorry i know there is a formula for the second part im just not sure how to use it in this instance

OpenStudy (anonymous):

Taylor series is this \[\sum_{n=0}^\infty\frac{f^n(0)}{n!}(x-a)^n\]

OpenStudy (anonymous):

using this series you have to compute n derivatives. usually your teacher will tell you oho many terms

OpenStudy (anonymous):

then after you're calculating the convergence of the series

OpenStudy (anonymous):

Error bound: | f^6 (x) | ≤ Max | f^6 (x) | = |- ( .3x)^6 | = ( .3x)^6 , x ≤ 1 => | f^6 (x) | ≤ ( .3 )^6 Max Since Error ≤ 1/ 10^k ( .3 )^6 / 6! ≤ 1/ 10^k -> k ≤ 5.99 ≈ 6